40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3099     Entry time: Tue Jun 22 20:07:08 2010
Author: Jenne 
Type: Update 
Category: 40m Upgrading 
Subject: First attempt at Tip Tilt hanging 

[Jenne, Steve, Nancy, Gopal]

We made an attempt at hanging some of the Tip Tilt eddy current dampers today. 

Photo 1 shows the 2 ECDs suspended.

Procedure:

(1) Loosen the #4-40 screws on the side of the ECDs, so the wire can be threaded through the clamps.

(2) Place the ECDs in the locator jigs (not shown), and the locator jigs in the backplane (removed from main TT structure), all laying flat on the table.

(3) Get a length of Tungsten wire (0.007 inch OD = 180um OD), wipe it with acetone, and cut it into 4 ~8cm long segments (long enough to go from the top of the backplane to the bottom).

(4) Thread a length of wire through the clamps on the ECDs, one length going through both ECDs' clamps.

(5) One person hold the wire taught, and straight, and as horizontal as possible, the other person tightened the clamping screws on the ECDs.

(6) Again holding the wire in place, one person put the clamps onto the backplane (the horizontal 'sticks' with 3 screws in them).

(7) The end. In the future, we'll also clip off extra pieces of wire.

When we held up the backplane to check out our handy work, it was clear that the bottom ECD was a much softer pendulum than the top one, since the top one has the wire held above and below, while the bottom one only has the wire held on the top.  I assume we'll trim the wire so that the upper ECD is only held on the top as well?

Lessons learned:

* This may be a 3 person job, or a 2 people who are good at multitasking job.  The wire needs to be held, the ECDs need to be held in place so they don't move during the screwing/clamping process, and the screws need to be tightened.

* Make sure to actually hold the wire taught. This didn't end up happening successfully for the leftmost wire in the photo, and the wire is a bit loose between the 2 ECDs.  This will need to be redone.

* We aren't sure that we have the correct screws for the clamps holding the wire to the backplane.  We only have 3/16" screws, and we aren't getting very many threads into the aluminum of the backplane.  Rana is ordering some 316 Stainless Steel (low magnetism) 1/4" #4-40 screws.  We're going for Stainless because Brass (the screws in the photo), while they passed their RGA scan, aren't really good for the vacuum.  And titanium is very expensive.  

The 2nd photo is of the magnet sticking out of the optic holder.  The hole that the magnet is sitting in has an aluminum piece ~2/3 of the way through.  A steel disk has been placed on one side, and the magnet on the other.  By doing this, we don't need to do any press-fitting (which was a concern whether or not the magnets could withstand that procedure), and we don't need to do any epoxying.  We'll have to wait until the ECDs are hung, and the optic holder suspended, to see whether or not the magnet is sticking out far enough to get to the ECDs. 

Attachment 1: 2_ECDs_small.jpg  114 kB  | Show | Show all
Attachment 2: MagnetStickingOutFar_small.jpg  73 kB  | Show | Show all
ELOG V3.1.3-